OpenStudy (punchthis22):

Ina shoots a large marble (Marble A, mass: 0.08 kg) at a smaller marble (Marble B, mass: 0.05 kg) that is sitting still. Marble A was initially moving at a velocity of 0.5 m/s, but after the collision it has a velocity of –0.1 m/s. What is the resulting velocity of marble B after the collision? Be sure to show your work for solving this problem along with the final answer.

OpenStudy (evoker):

Use conservation of momentum so mv+mv=mv+mv

OpenStudy (osprey):

Using conservation of linear momentum with the numbers given, I can work out a resulting velocity for v. BUT ... To check it, I’ve tried to apply conservation of kinetic energy, or at least COMPARE the kinetic energies before and after collision. See enclosed powerpoint http://perendis.webs.com

OpenStudy (sharmagk83):

Here we will use law of conservation of momentum i.e. Total momentum before collision = total momentum after collision \[m_{1} u_{1} +m_{2} u_{2}= m_{1}v_{1}+m_{2}v_{2}\] here \[m_{1} = 0.08 kg\] , \[u_{1} = 0.5 m/s\], \[m_{2} = 0.05 kg\], \[u_{2} = 0\], \[v_{1} = -1 m/s\] Now \[0.08 kg \times 0.5 m/s +0.05 kg \times0 m/s= 0.08 kg \times (-0.1m/s)+ 0.05\times(v_{2})\] \[v_{2} = \frac{ 0.08\times0.5 +0.08\times0.1}{ 0.05 }m/s\] \[v_{2} = 0.816 m/s\]